# Estimating Heterogeneous Causal Effects of High-Dimensional Treatments:

### v4

@@ 3.4 Sparsity-induced Prior
We encourage such fusion by applying a structured sparsity approach of Goplerud (2021) that generalizes the group and overlapping group LASSO (e.g., Yuan and Lin, 2006; Yan and Bien, 2017) while allowing positive semi-definite constraint matrices. For computational tractability, we use $\ell_{2}$ norm instead of the $\ell_{\infty}$ norm, which is used in GASH-ANOVA (Post and Bondell, 2013). An additional benefit of the use of regularization is that it gives us some protection against finding spurious relations (see Gelman, Hill and Yajima, 2012).
@@
We generalize this formulation to an arbitrary number of factors and factor levels. For each factor that contains $L_{j}$ levels, we have $\binom{L_{j}}{2}$ penalty matrices to encourage pairwise fusion. Imposing additional constraints is a simple extension; for example, for ordered factors, one might use penalties that penalize the differences between adjacent levels (e.g. $l$ and $l+1$). Let $G=\sum_{j=1}^{J}\binom{L_{j}}{2}$ represent the total number of penalty matrices. For $g=1,2,\ldots,G$, we use $\bm{F}_{g}$ to denote a penalty matrix such that $\sqrt{\bm{\beta}^{\top}\bm{F}_{g}\bm{\beta}}$ is equivalent to the $\ell_{2}$ norm on the vector of differences between all main effects and interactions containing a main effect. We note that $\{\bm{F}_{g}\}_{g=1}^{G}$ is not directly chosen but rather are determined by factors in the experiment ($J$, $L_{j}$, whether $j$ is ordered or unordered) and the included interactions (as well as the use of “latent overlapping groups”; see Appendix G.4).
@@
$p\left(\bm{\beta}_{k}\mid\{\bm{\phi}_{k}\}_{k=2}^{K}\right)\ \propto\ \left(\lambda{\bar{\pi}_{k}^{\gamma}}\right)^{m}\exp\left(-\lambda\bar{\pi}_{k}^{\gamma}\sum_{g=1}^{G}\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}\right),$ (7)
where $\bar{\pi}_{k}=\sum_{i=1}^{N}\pi_{k}(\bm{X}_{i})/N$ and $m=\text{rank}\left([\bm{F}_{1},\cdots,\bm{F}_{G}]\right)$. We follow existing work in allowing the penalty on the treatment effects $\bm{\beta}_{k}$ to be scaled by the group-membership size $\bar{\pi}_{k}$ when $\gamma=1$ (Khalili and Chen, 2007; Städler, Bühlmann and Van De Geer, 2010). On the other hand, when $\gamma=0$ the $\bar{\pi}_{k}$ disappears, implying no use of the $\bm{X}_{i}$ in the prior. We note that the prior on $p(\bm{\beta}\mid\{\bm{\phi}_{k}\}_{k=2}^{K})$ is guaranteed to be proper when all pairwise fusions are encouraged by $\{\bm{F}_{g}\}_{g=1}^{G}$, although in other circumstances it may be improper (Goplerud, 2021). Appendix E provides additional details. Following Zahid and Tutz (2013), we use a normal prior distribution for the coefficients for the moderators.
@@
$p(\{\bm{\phi}_{k}\}_{k=2}^{K})\ \propto\ \exp\left(-\frac{\sigma^{2}_{\phi}}{2}\sum_{l=1}^{p_{x}}[\bm{\phi}_{2l},\cdots,\bm{\phi}_{Kl}]^{\top}\bm{\Sigma}_{\phi}[\bm{\phi}_{2l},\cdots,\bm{\phi}_{Kl}]\right),$ (8)
@@ 3.5 Estimation and Inference
We fit our model by finding a maximum of the log-posterior using an extension of the Expectation-Maximization (EM; Dempster, Laird and Rubin 1977) algorithm known as the Alternating Expectation-Conditional Maximization (AECM; Meng and van Dyk 1997) algorithm. Equation (9) defines our (observed) log-posterior using the terms defined in Equations (3), (7), and (8), where we collect all model parameters as $\bm{\theta}$:
$\begin{split}\log p\left(\bm{\theta}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N}\right)\ =\ &\sum_{i=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right]+\\ &\sum_{k=1}^{K}\log p(\bm{\beta}_{k}\mid\{\bm{\phi}\}_{k=2}^{K})+\log p(\{\bm{\phi}_{k}\}_{k=2}^{K})+\text{const}.\end{split}$ (9)
For now, we assume the value of regularization parameter $\lambda$ is fixed, although we discuss this issue in Section 3.6. The linear constraints on $\bm{\beta}_{k}$ given in Equation (6) still hold but are suppressed for notational simplicity.
Appendix F provides a full derivation of our AECM algorithm; each iteration involves two cycles where the data augmentation scheme enables iterative updating of the treatment effect parameters $\bm{\beta}$ and moderators $\bm{\phi}$. After augmenting with missing data, the update for $\bm{\beta}$ can be done using ridge regression; Appendix F.1 addresses the linear constraints imposed by $\bm{C}^{T}\bm{\beta}_{k}=\bm{0}$. The update for $\bm{\phi}$ can be performed using a modified version of a multinomial logistic regression based on a standard optimizer (e.g., L-BFGS) (see Appendix F.2).
@@ 3.6 Additional Considerations
Since fitting the proposed model is computationally expensive, we use the Bayesian Information Criteria (BIC), rather than cross validation, to select the value of the regularization parameter $\lambda$ (Khalili and Chen, 2007; Khalili, 2010; Chamroukhi and Huynh, 2019). Appendix F.3 presents our degrees-of-freedom estimator and explains how we tune $\lambda$ using Bayesian model-based optimization. Appendix F.4 discusses additional details of our EM algorithm including initialization and techniques to accelerate convergence.
@@
We extend the above model and estimation algorithm to accommodate common features of conjoint analysis: (1) repeated observations for each individual respondent (Appendix G.1), (2) a forced choice conjoint design (Appendix G.2), and (3) standardization weights for factors with different numbers of levels $L_{j}$ (Appendix G.3). Lastly, our experience suggests that the proposed penalty function, which consists of overlapping groups, often finds highly sparse solutions. Appendix G.4 details the integration of the latent overlapping group formulation of Yan and Bien (2017) into our framework to address this issue.
@@
Once the model parameters are estimated, we can compute quantities of interest such as the AMCEs, defined in Equation (1). We do this separately for each group, such that $\delta_{jk}(l,l^{\prime})$ is the AMCE for factor $j$, changing from level $l^{\prime}$ to $l$ in group $k$. Our estimator is the average of the estimated difference in predicted responses when changing from level $l^{\prime}$ to $l$ of factor $j$, where the average is taken over the empirical distribution of the assignment on the other factors. This estimation is described in more detail in Appendix H under various settings. We can use the empirical distribution here because treatment is randomly assigned.
@@
To quantify the uncertainty of the parameter estimates, we rely on a quadratic approximation to the log posterior distribution. To ensure its differentiability, we follow a standard approach in the regularized regression literature (e.g., Fan and Li, 2001) and fuse pairwise factor levels that are sufficiently close together. Appendix I describes this process, deriving the Hessian of the log-posterior using Louis (1982)’s method and then using the delta method for inference on other of quantities of interest, e.g., the AMCE.
@@
Finally, in principle, our framework does not assume a “correct” data generating process. The choice of number of groups $K$ should depend on the desired granularity of discovered heterogeneity, with more groups leading to finer levels of heterogeneity. Similarly, the choice of moderators should reflect the researcher’s substantive interests. Appendix J.3 shows performance of our method across different values of $K$ and different specifications of the moderators when the true data generating process is a mixture model. As expected, the bias of AMCE is not affected by changing the specification of these parameters. However, there are some impacts on the estimation of conditional effects in terms of precision.
@@ 4 Simulations
In all cases, we assume that the data generating process follows a mixture of experts model with three groups ($K=3$). We calibrate the true $\bm{\beta}_{k}$ such that the implied average marginal component effects (AMCE) are comparable in magnitude to the empirical effects presented in Section 5. We use a set of five correlated continuous moderators and an intercept to again mimic a realistic empirical setting and choose $\{\bm{\phi}_{k}\}_{k=2}^{3}$ to relatively clearly separate respondents into different groups. Appendix J presents complete description of the simulation settings and the true parameter values used for the $\bm{\beta}_{k}$ and marginal effects.
@@
Fig 2 summarizes our results (see Appendix J.2 for the results regarding the estimated coefficients $\bm{\beta}_{k}$). The left panel illustrates a high correlation between the estimated effects and their true values ($\rho=0.995$ for smaller sample size; $\rho=0.999$ for larger sample size). While the performance overall is reasonably strong, we see that even when the dataset is large there is some degree of attenuation bias due to shrinkage.
@@
Appendix J.2 explores one way to address the limitations of the default estimator by exploring sample splitting and refitting the model given the estimated sparsity pattern (i.e., which levels are fused together) and moderator effects ($\{\bm{\phi}_{k}\}_{k=2}^{K}$) on half of the data. This results in smaller bias and improved coverage at both sample sizes.
Appendix J.3 explores how when the true data generating process is a mixture model, the “wrong” choice of $K$, e.g. $K\in\{1,2,4\}$, as well as not using moderators (i.e., $\bm{X}_{i}=1$) or using moderators in a different specification than the true model impacts our results. In both settings, there is limited impact in terms of bias in terms of estimating the AMCE, although both types of misspecification incur a penalty in terms of root mean-squared error.
@@ 5.1 Data and Model
To account for randomization restrictions, we include interactions between country of origin and reason for applying as well as between job and education. To test relevant theories, we include additional interactions between country of origin and job as well as country of origin and education. We focus on only these two factor interactions to keep the model interpretable. We then appeal to the hierarchy and sparsity principles (Wu and Hamada, 2021), which imply that lower-order effects are expected to be more significant than higher-order effects and we should expect a relatively small number of nonzero effects. With this linear predictor formulation, the estimation and inference proceed as explained in Section 3.
We conduct two analyses, one with two groups and the other with three groups. These two models perform equally well in terms of out-of-sample classification, a data-driven measure that can be used to choose the number of groups. Using more than three groups does not give improved substantive insights and provides little improvement in model performance. As noted previously, each analysis optimizes the BIC to calibrate the amount of regularization and employs standardization weights to account for factors with different number of levels (see Appendices F.3 and G.3, respectively, for details). We treat education and job experience as ordered factors and only penalize the differences between adjacent levels.
We report our findings using only the full data estimates, i.e., without the sample splitting explored in Appendix J.2. Initial experiments found that the results were somewhat sensitive to specific folds chosen, and thus we report only the full data results in the main text. Appendix K illustrates the distribution of estimates across twenty different sample splits.
@@ 5.2 Estimated Heterogeneity
Figure 3: Estimated average marginal component effects using a two-group (left) and three-group (right) analysis. The point estimates and 95% Bayesian credible intervals are shown. A solid circle represents the baseline level of each factor. Number after colon give average posterior predictive probability for each group.
@@
For both analyses, the respondents in Group 1 do not care much about immigrant’s country of origin. Instead, they place a greater emphasis on education and reason for immigration when compared to those in the other groups. For the three-group analysis, Groups 1 and 2 together correspond roughly to Group 1 of the two-group analysis. In fact, about 81% of the respondents who belong to Group 1 of the two-group analysis are the members of either Group 1 or 2 in the three-group analysis using a weighted average of their estimated group membership posterior predictive probabilities. The differences between Groups 1 and 2 in the three-group analysis are substantively small, but those in Group 2 appear to place more emphasis on education and prior entry without legal authorization. Those in Group 1, on the other hand, give a slight benefit to immigrants whose reason for immigration is persecution.
@@
The comparison of AMCEs across subgroups can be misleading as they depend on the choice of baseline category (Leeper, Hobolt and Tilley, 2020). Appendix K presents an alternative quantity that avoids issues of baseline dependency (marginal means; Leeper, Hobolt and Tilley 2020). The results are generally similar to AMCEs shown above.

### v5

@@ 3.4 Sparsity-inducing Prior
We encourage such fusion by applying a structured sparsity approach of Goplerud (2021) that generalizes the group and overlapping group LASSO (e.g., Yuan and Lin, 2006; Yan and Bien, 2017) while allowing positive semi-definite penalty matrices. For computational tractability, we use $\ell_{2}$ norm instead of the $\ell_{\infty}$ norm, which is used in GASH-ANOVA (Post and Bondell, 2013). An additional benefit of the use of regularization is that it gives us some protection against finding spurious relations (see Gelman, Hill and Yajima, 2012).
@@
where $\bm{F}_{1},\bm{F}_{2},\bm{F}_{3}$ are appropriate positive semi-definite matrices to encourage the fusion of the pairs of levels in factor one and $\bm{F}_{4}$ encourages the fusion of the two levels in factor two, and
@@
We generalize this formulation to an arbitrary number of factors and factor levels. For each factor that contains $L_{j}$ levels, we have $\binom{L_{j}}{2}$ penalty matrices to encourage pairwise fusion. Imposing additional constraints is a simple extension; for example, for ordered factors, one might use penalties that penalize the differences between adjacent levels (e.g. $l$ and $l+1$). Let $G=\sum_{j=1}^{J}\binom{L_{j}}{2}$ represent the total number of penalty matrices. For $g=1,2,\ldots,G$, we use $\bm{F}_{g}$ to denote a penalty matrix such that $\sqrt{\bm{\beta}^{\top}\bm{F}_{g}\bm{\beta}}$ is equivalent to the $\ell_{2}$ norm on the vector of differences between all main effects and interactions containing a main effect. We note that $\{\bm{F}_{g}\}_{g=1}^{G}$ is not directly chosen but rather are determined by factors in the experiment ($J$, $L_{j}$, whether $j$ is ordered or unordered) and the included interactions (as well as the use of “latent overlapping groups”; see Section H.4 of the Supplementary Material.
@@
$p\left(\bm{\beta}_{k}\mid\{\bm{\phi}_{k}\}_{k=2}^{K}\right)\ \propto\ \left(\lambda{\bar{\pi}_{k}^{\gamma}}\right)^{m}\exp\left(-\lambda\bar{\pi}_{k}^{\gamma}\sum_{g=1}^{G}\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}\right),$ (3.8)
where $\bar{\pi}_{k}=\sum_{i=1}^{N}\pi_{k}(\bm{X}_{i})/N$ and $m=\text{rank}\left([\bm{F}_{1},\cdots,\bm{F}_{G}]\right)$. We follow existing work in allowing the penalty on the treatment effects $\bm{\beta}_{k}$ to be scaled by the group-membership size $\bar{\pi}_{k}$ when $\gamma=1$ (Khalili and Chen, 2007; Städler, Bühlmann and Van De Geer, 2010). On the other hand, when $\gamma=0$ the $\bar{\pi}_{k}$ disappears, implying no use of the $\bm{X}_{i}$ in the prior. We note that the prior on $p(\bm{\beta}\mid\{\bm{\phi}_{k}\}_{k=2}^{K})$ is guaranteed to be proper when all pairwise fusions are encouraged by $\{\bm{F}_{g}\}_{g=1}^{G}$, although in other circumstances it may be improper (Goplerud, 2021). Section F of the Supplementary Material provides additional details. Following Zahid and Tutz (2013), we use a normal prior distribution for the coefficients for the moderators.
@@
$p(\{\bm{\phi}_{k}\}_{k=2}^{K})\ \propto\ \exp\left(-\frac{\sigma^{2}_{\phi}}{2}\sum_{l=1}^{p_{x}}[\bm{\phi}_{2l},\cdots,\bm{\phi}_{Kl}]^{\top}\bm{\Sigma}_{\phi}[\bm{\phi}_{2l},\cdots,\bm{\phi}_{Kl}]\right),$ (3.9)
@@ 3.5 Estimation and Inference
We fit our model by finding a maximum of the log-posterior using an extension of the Expectation-Maximization (EM; Dempster, Laird and Rubin 1977) algorithm known as the Alternating Expectation-Conditional Maximization (AECM; Meng and van Dyk 1997) algorithm. Equation (3.10) defines our (observed) log-posterior using the terms defined in Equations (3.3), (3.8), and (3.9), where we collect all model parameters as $\bm{\theta}$:
$\begin{split}\log p\left(\bm{\theta}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N}\right)\ =\ &\sum_{i=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right]+\\ &\sum_{k=1}^{K}\log p(\bm{\beta}_{k}\mid\{\bm{\phi}\}_{k=2}^{K})+\log p(\{\bm{\phi}_{k}\}_{k=2}^{K})+\text{const}.\end{split}$ (3.10)
For now, we assume the value of regularization parameter $\lambda$ is fixed, although we discuss this issue in Section 3.6. The linear constraints on $\bm{\beta}_{k}$ given in Equation (3.7) still hold but are suppressed for notational simplicity.
Section G of the Supplementary Material provides a full derivation of our AECM algorithm; each iteration involves two cycles where the data augmentation scheme enables iterative updating of the treatment effect parameters $\bm{\beta}$ and moderators $\bm{\phi}$. After augmenting with missing data, the update for $\bm{\beta}$ can be done using ridge regression; Section G.1 addresses the linear constraints imposed by $\bm{C}^{T}\bm{\beta}_{k}=\bm{0}$. The update for $\bm{\phi}$ can be performed using a modified version of a multinomial logistic regression based on a standard optimizer (e.g., L-BFGS) (see Section G.2).
@@ 3.6 Additional Considerations
Since fitting the proposed model is computationally expensive, we use the Bayesian Information Criteria (BIC), rather than cross validation, to select the value of the regularization parameter $\lambda$ (Khalili and Chen, 2007; Khalili, 2010; Chamroukhi and Huynh, 2019). Section G.4 of the Supplementary Material presents our degrees-of-freedom estimator and explains how we tune $\lambda$ using Bayesian model-based optimization. Section G.5 discusses additional details of our EM algorithm including initialization and techniques to accelerate convergence.
@@
We extend the above model and estimation algorithm to accommodate common features of conjoint analysis: (1) repeated observations for each individual respondent (Section H.1 of the Supplementary Material), (2) a forced choice conjoint design (Section H.2), and (3) standardization weights for factors with different numbers of levels $L_{j}$ (Section H.3). Lastly, our experience suggests that the proposed penalty function, which consists of overlapping groups, often finds highly sparse solutions. Section H.4 details the integration of the latent overlapping group formulation of Yan and Bien (2017) into our framework to address this issue.
@@
Once the model parameters are estimated, we can compute quantities of interest such as the AMCEs, defined in Equation (3.1). We do this separately for each group, such that $\delta_{jk}(l,l^{\prime})$ is the AMCE for factor $j$, changing from level $l^{\prime}$ to $l$ in group $k$. Our estimator is the average of the estimated difference in predicted responses when changing from level $l^{\prime}$ to $l$ of factor $j$, where the average is taken over the empirical distribution of the assignment on the other factors. This estimation is described in more detail in Section I of the Supplementary Material under various settings. We can use the empirical distribution here because treatment is randomly assigned.
@@
To quantify the uncertainty of the parameter estimates, we rely on a quadratic approximation to the log posterior distribution. To ensure its differentiability, we follow a standard approach in the regularized regression literature (e.g., Fan and Li, 2001) and fuse pairwise factor levels that are sufficiently close together. Section J of the Supplementary Material describes this process, deriving the Hessian of the log-posterior using Louis (1982)’s method and then using the delta method for inference on other of quantities of interest, e.g., the AMCE.
@@
Finally, in principle, our framework does not assume a “correct” data generating process. The choice of number of groups $K$ should depend on the desired granularity of discovered heterogeneity, with more groups leading to finer levels of heterogeneity. Similarly, the choice of moderators should reflect the researcher’s substantive interests. Section K.3 of the Supplementary Material shows performance of our method across different values of $K$ and different specifications of the moderators when the true data generating process is a mixture model. As expected, the bias of AMCE is not affected by changing the specification of these parameters. However, there are some impacts on the estimation of conditional effects in terms of precision.
Common data-driven approaches for choosing $K$ include use of an information criterion such as the BIC; however, while we find that these approaches work well under simulation settings (see Section K.3.1 for demonstration), they can perform poorly in practice (see Section L), especially when the component densities are mis-specified or not especially well separated (Celeux, Frühwirth-Schnatter and Robert, 2019). Thus, even if a data-driven heuristic is used as a guide for choosing $K$, we suggest comparing different $K$ as illustrated in Section 5.
@@ 4 Simulations
In all cases, we assume that the data generating process follows a mixture of experts model with three groups ($K=3$). We calibrate the true $\bm{\beta}_{k}$ such that the implied average marginal component effects (AMCE) are comparable in magnitude to the empirical effects presented in Section 5. We use a set of five correlated continuous moderators and an intercept to again mimic a realistic empirical setting and choose $\{\bm{\phi}_{k}\}_{k=2}^{3}$ to relatively clearly separate respondents into different groups. Section K of the Supplementary Material presents complete description of the simulation settings and the true parameter values used for the $\bm{\beta}_{k}$ and marginal effects.
@@
Fig 2 summarizes our results (see Section K.2 for the results regarding the estimated coefficients $\bm{\beta}_{k}$). The left panel illustrates a high correlation between the estimated effects and their true values ($\rho=0.995$ for smaller sample size; $\rho=0.999$ for larger sample size). While the performance overall is reasonably strong, we see that even when the dataset is large there is some degree of attenuation bias due to shrinkage.
@@
Section K.2 explores one way to address the limitations of the default estimator by exploring sample splitting and refitting the model given the estimated sparsity pattern (i.e., which levels are fused together) and moderator effects ($\{\bm{\phi}_{k}\}_{k=2}^{K}$) on half of the data. This results in smaller bias and improved coverage at both sample sizes.
Section K.3 explores how when the true data generating process is a mixture model, the “wrong” choice of $K$, e.g. $K\in\{1,2,4\}$, as well as not using moderators (i.e., $\bm{X}_{i}=1$) or using moderators in a different specification than the true model impacts our results. In both settings, there is limited impact in terms of bias in terms of estimating the AMCE, although both types of misspecification incur a penalty in terms of root mean-squared error.
@@ 5.1 Data and Model
To account for randomization restrictions, we include interactions between country of origin and reason for applying as well as between job and education. To test relevant theories, we include additional interactions between country of origin and job as well as country of origin and education. These interaction effects proved to be very small in magnitude (see Section L of the Supplementary Material). Thus, we do not explore higher order interactions given the commonly adopted principle of hierarchy and sparsity (Wu and Hamada, 2021), which implies that lower-order effects are expected to be more significant than higher-order effects and we should expect an even smaller number of nonzero effects. With this linear predictor formulation, the estimation and inference proceed as explained in Section 3.
We conduct two analyses, one with two groups and the other with three groups. These two models perform equally well in terms of out-of-sample classification, a data-driven measure that can be used to choose the number of groups. Using more than three groups does not give improved substantive insights and provides little improvement in model performance. As noted previously, each analysis optimizes the BIC to calibrate the amount of regularization and employs standardization weights to account for factors with different number of levels (see Sections G.4 and H.3 of the Supplementary Material, respectively, for details). We treat education and job experience as ordered factors and only penalize the differences between adjacent levels.
We report our findings using only the full data estimates, i.e., without the sample splitting explored in Section K.2. Initial experiments found that the results were somewhat sensitive to specific folds chosen, and thus we report only the full data results in the main text. Section L illustrates the distribution of estimates across twenty different sample splits.
@@ 5.2 Estimated Heterogeneity
Figure 3: Estimated average marginal component effects using a two-group (left) and three-group (right) analysis. The point estimates and 95% Bayesian credible intervals are shown. A solid circle represents the baseline level of each factor. Numbers after colons give average posterior predictive probabilities for each group.
@@
For both analyses, the respondents in Group 1 do not care much about immigrant’s country of origin. Instead, they place a greater emphasis on education and reason for immigration when compared to those in the other groups. While the differences between Groups 1 and 2 in the three-group analysis are generally substantively small, those in Group 2 appear to place more emphasis on education and prior entry without legal authorization. Those in Group 1, on the other hand, give a slight benefit to immigrants whose reason for immigration is persecution.
Figure 4: Ternary plot of the joint posterior predictive probability of belonging to each group in the three-group analysis (three axes) where the color of each dot represents the posterior predictive probability of belonging to Group 1 under the two-group analysis.
Indeed, for the three-group analysis, Groups 1 and 2 together correspond roughly to Group 1 of the two-group analysis. In fact, about 81% of the respondents who belong to Group 1 of the two-group analysis are the members of either Group 1 or 2 in the three-group analysis, using a weighted average of their estimated group membership posterior predictive probabilities.
Fig 4 visualizes these posterior predictive probabilities of group membership under the three-group analysis with each dot colored by the posterior predictive probability of belonging to Group 1 under the two-group analysis. According to this ternary plot, those observations that are likely to be part of Group 1 under the two-group analysis (i.e., red dots) are likely to be split between Groups 2 and 3 under the three group analysis. In contrast, those who have a high probability of belonging to Group 2 under the two-group analysis (i.e., blue dots) tend to be part of Group 3.
@@
The comparison of AMCEs across subgroups can be misleading as they depend on the choice of baseline category (Leeper, Hobolt and Tilley, 2020). Section L of the Supplementary Material presents an alternative quantity that avoids issues of baseline dependency (marginal means; Leeper, Hobolt and Tilley 2020). The results are generally similar to AMCEs shown above.